OpenStudy (anonymous):

256x^2-100 Please show me step by step. I really want to make sure i got this down.Thanks

OpenStudy (campbell_st):

well this is a quadratic with only 2 terms and no term involving x so its the difference of 2 squares (a + b)(a-b) = a^2 - b^2 256x^2 = (16x)^2 100 = (10)^2 so the factorisation is (16x - 10)(16x + 10)

OpenStudy (anonymous):

thanks very much